# MVTamperBench: Evaluating Robustness of Vision-Language Models

### v4

@@ Abstract
Multimodal Large Language Models (MLLMs), also known as Large Multi-modal Models (LMMs), are recent advancement of Vision-Language Models (VLMs), that have driven major advances in video understanding, yet their vulnerability to adversarial tampering and manipulations remains underexplored. To address this gap, we introduce MVTamperBench, a benchmark that systematically evaluates MLLM robustness against five prevalent tampering techniques: rotation, masking, substitution, repetition, and dropping. Built from 3.4K original videos—expanded to over 17K tampered clips spanning 19 video tasks.
MVTamperBench challenges models to detect manipulations in spatial and temporal coherence. We evaluate 45 recent MLLMs from 15+ model families, revealing substantial variability in resilience across tampering types and showing that larger parameter counts do not necessarily guarantee robustness. MVTamperBench sets a new benchmark for developing tamper-resilient MLLM in safety-critical applications, including detecting clickbait, preventing harmful content distribution, and enforcing policies on media platforms. We release all code [^1] https://amitbcp.github.io/MVTamperBench/ and data [^2] https://huggingface.co/datasets/Srikant86/MVTamperBench to foster open research in trustworthy video understanding.
@@
Amit Agarwal1*, Srikant Panda2*, Angeline Charles3*, Bhargava Kumar4, Hitesh Patel5, Priyaranjan Pattnayak6, Taki Hasan Rafi7, Tejaswini Kumar 4, Dong-Kyu Chae7+ 1Liverpool John Moores University 2Birla Institute of Technology 3Christ University 4Columbia University 5New York University 6University of Washington 7Hanyang University *Equal Contribution +Leadership Correspondence: amit.pinaki@gmail.com, dongkyu@hanyang.ac.kr
@@ 1 Introduction
Multimodal Large Language Models (MLLMs) have driven significant advancements in video understanding, enabling applications across domains such as surveillance, healthcare, and autonomous systems. However, their increasing adoption on high-usage platforms (e.g., Instagram, Facebook) has exposed critical vulnerabilities. Tampered videos are frequently weaponized to evade platform policies, spread harmful content, and propagate clickbait, posing significant challenges for content moderation and policy enforcement Kingra et al. (2023); Times of India (2024). These vulnerabilities underscore the urgent need to enhance the robustness of MLLMs against real-world manipulations.
While adversarial robustness has been extensively studied in static image contexts, video tampering presents unique challenges. Techniques such as frame dropping, masking, repetition, substitution, and rotation disrupt both spatial and temporal coherence, often causing severe model failures. Existing adversarial approaches, including black-box Jiang et al. (2019) and cross-modal Wei et al. (2021) attacks, primarily address isolated scenarios rather than systematically evaluating diverse tampering types.
Existing multimodal benchmarks Pattnayak et al. (2024); AGARWAL (2021)—such as MMBench-Video Fang et al. (2024), BLINK Fu et al. (2025), and Video-MME Fu et al. (2024)—focus on multimodal comprehension and temporal reasoning but overlook adversarial robustness. For example, MMBench-Video evaluates cross-modal alignment without adversarial testing; BLINK targets long-form temporal reasoning without addressing tampering impacts; and Video-MME, while effective for vision-language alignment, omits tampering-specific tasks. This leaves a critical gap in systematically assessing how MLLMs withstand real-world manipulations.
To bridge this gap, we introduce MVTamperBench, a benchmark specifically designed to evaluate MLLM robustness against five prevalent tampering techniques. Built from 3.4K original videos—expanded into over 17K tampered clips spanning 19 video tasks—MVTamperBench challenges models to detect manipulations by distrupting temporal and spatial coherence in a video.
@@
  • Through experiments on 45 MLLMs across 15+ families, we identify critical vulnerabilities across MLLM families, without any correlation between model size and performance.
By focusing on tampering resilience, MVTamperBench provides a critical tool for improving MLLM robustness. It enables the development of tamper-resilient models applicable to real-world challenges such as clickbait detection, content moderation, and policy enforcement, advancing adversarial robustness in video understanding.
@@ 2 Related Work
The development of benchmarks for MLLMs has significantly advanced the evaluation of image and video understanding tasks, encompassing spatial reasoning, temporal comprehension, object detection, and common sense inference. However, the robustness of MLLMs to adversarial manipulations, particularly video tampering, remains underexplored. This section reviews existing benchmarks, categorizing them into image-based and video-based evaluations, and critically analyzes their contributions and shared limitations.
@@ 2.1 Image-Based Understanding
Benchmarks like BLINK Fu et al. (2025) and MuirBench Wang et al. (2024a) focus on evaluating static visual reasoning. BLINK tests foundational tasks such as depth estimation, forensic detection, and visual correspondence, providing a diverse set of challenges for spatial reasoning. Similarly, MuirBench extends this evaluation to multi-image tasks, including action recognition and geographic reasoning, by synthesizing information from diverse sources. While these benchmarks have advanced static image understanding, their reliance on single or multiple still images excludes temporal dynamics and limits their applicability to scenarios involving sequential manipulations, such as those found in video content.
@@ 2.2 Video-Based Understanding
2.3 Identified Gaps and Motivation
Despite their contributions to advancing MLLM evaluation, existing benchmarks largely focus on performance under ideal conditions, neglecting robustness to tampering or adversarial effects. Techniques such as frame substitution, masking, repetition, dropping, and rotation disrupt temporal coherence and pose unique challenges for multimodal models. The absence of systematic evaluations of these tampering techniques highlights a critical gap in ensuring model reliability for real-world applications like forensic analysis, media verification, and misinformation detection.
Figure 1: Illustration of Video Frame Tampering Techniques. Each row shows a specific tamper type applied to a video snippet, highlighting the respective impact on temporal/spatial coherence.
To address these challenges, we introduce MVTamperBench, a comprehensive benchmark designed to systematically evaluate MLLMs robustness against video tampering effects. By introducing diverse manipulations, MVTamperBench broadens the evaluation landscape, enabling a deeper understanding of model strengths and vulnerabilities under adversarial scenarios.
@@ 3 MVTamperBench
We introduce MVTamperBench, our benchmark specifically designed to evaluate the robustness of MLLMs against diverse video tampering techniques. In the following subsections, we detail its construction, design choices, and key features.
@@ 3.1 Benchmark Construction
To evaluate MLLM robustness under adversarial conditions, we apply distinct tampering methods—Dropping, Masking, Substitution, Repetition, and Rotation—to the 3,487 original MVBench Li et al. (2024e) videos (excluding NTU dataset due to licensing), resulting in a total of 17,435 tampered clips. These manipulations target both spatial and temporal coherence, simulating common real-world tampering scenarios such as deliberate frame editing or slicing from unrelated content.
@@ 3.2 Tampering Techniques
Dropping: Removes a 1-second segment, creating temporal discontinuity.
Masking: Overlays a black rectangle on a 1-second segment, simulating visual data loss.
Rotation: Rotates a 1-second segment by 180 degrees, introducing spatial distortion.
Substitution: Replaces a 1-second segment with a pre-selected clip from another video, disrupting temporal and contextual flow.
@@
By applying these effects uniformly across all videos, MVTamperBench ensures consistent and comparable evaluation.
@@ 3.3 Design & Implementation
Tampering Duration (1 Second). We fix tampering to a 1-second segment to align with reports of minimal yet impactful real-world tampering, e.g., short edits in social media or subtle modifications in surveillance feeds Alicia (2024). Preliminary experiments revealed that using less than 1 second could be overlooked by certain model sampling mechanisms, whereas longer tampering (>3s) often resembled normal scene transitions, reducing adversarial impact.
Tampering Location (Middle). All manipulations occur at the video’s midpoint to disrupt central content. Pilot tests showed that tampering near the start or end risked mimicking scene cuts or information loss, making detection less indicative of genuine adversarial robustness.
Substitution Source. For Substitution, the 1-second clip is randomly chosen from a consistent pool of different videos within MVBench. This ensures uniform difficulty across all samples, preventing confounds from domain shifts or overly simplistic substitutes.
Modular & Scalable Framework. Each technique is encapsulated in a reusable class, facilitating custom parameterization (e.g., tampering duration, location, intensity). Our open-source code integrates seamlessly with VLMEvalKit Duan et al. (2024), promoting reproducible experiments and easy extension to other datasets or tasks.
@@ 3.4 Dataset Scope and Statistics
All 3,487 MVBench videos undergo each of the five tampering types, producing 17,435 tampered clips. Figure 2 illustrates the distribution of their durations, capturing diverse scenarios from short (3-5s) to extended (>20s) sequences. By uniformly applying manipulations, MVTamperBench ensures comparability across different MLLM architectures and training regimes.
Figure 2: The dataset spans a broad range of durations, reflecting varied real-world conditions.
@@ 3.5 Summary
By enforcing consistent parameters (1s duration, midpoint placement) and systematically applying five tampering methods, MVTamperBench offers a controlled yet flexible platform for evaluating tampering resilience. Our design can be easily extended with additional manipulations (e.g., noise injection, partial masking) or integrated into domain-specific contexts like surveillance feeds analysis or clickbait detection. We further compare MVTamperBench with other benchmarks, details on source of videos and tasks in Appendix 7.1
@@
4 Experimentation and Results
This section outlines the experimental setup, results, and analysis of 45 models evaluated on the benchmark, highlighting their strengths, weaknesses, and actionable insights. We provide further overview of the evaluated MLLMs in Appendix 7.3.
Evaluation Protocol Each model is tasked with identifying whether a video has been tampered with or not, simulating real-world scenarios where detecting manipulations is critical. For every video, the model is presented with the following prompt:
@@ 4 Experimentation and Results
The dataset comprises both tampered and untampered videos. For each tampered video, the corresponding untampered video is included, ensuring a balanced distribution of tampered and untampered samples. Models must select one of the two options (Yes or No), and their predictions are compared against ground truth labels to determine correctness. This task is repeated for all tampering types, enabling a comprehensive assessment of model robustness across diverse manipulations.
Metrics The primary evaluation metric is the F1 Score, chosen for its ability to balance precision and recall, particularly in scenarios where misclassifications (false positives and false negatives) can significantly impact robustness evaluation. This is particularly important in our setup, where models must not only detect manipulations but also avoid false positives on non-tampered videos. To capture performance across all tampering types, we compute individual F1 scores for each tampering type. The F1 (overall) score is then calculated as the macro average of these individual F1 scores, ensuring all tampering types contribute equally to the final evaluation. This approach effectively highlights per-tamper-type strengths and weaknesses while providing a holistic measure of overall model performance.
@@
Figure 5: Correlation between model size and F1 (overall), with a Pearson value of 0.05.
@@ 4.1 Results and Analysis
We evaluate the performance of 45 MLLMs using F1 (overall) scores and individual F1 scores across five tampering types. The distribution of F1 (overall) scores (Figure 4) reveals significant variability in model robustness, with several models struggling to detect tampering (F1 < 0.2), while a few high-performing models achieve F1 > 0.8.
We observe no correlation between model size and tampering detection performance (Pearson correlation = 0.05, Figure 5). This highlights that architectural differences and training techniques, rather than parameter count, contribute more significantly to robustness against tampering.
Adaption to Tampering Techniques
Figure 3 showcases model-specific adaptability across different tampering types. Models with F1 (overall) > 0.8 are consistently robust across all tampering types, while those with F1 < 0.2 perform slightly better on Masking, which relies more on spatial reasoning. Models in between generally struggle with temporal disruptions like Dropping and Substitution, reflecting challenges in temporal coherence.
4.1.1 Models Analysis Based on Performance
Figure 6 categorizes MLLMs into low-, moderate-, and high -performing groups based on their F1-score distribution. We define the boundaries using the 0.25 quantile (F1 = 0.071) and 0.75 quantile (F1 = 0.846). To round-off the quantiles thresholds to the nearest integer, making the low-performing models (F1 (overall) < 0.01) and high-performing models (F1 (overall) > 0.8).
Figure 7 illustrates the average F1 scores for each category, revealing stark differences between groups. High-performing models maintain consistent performance across all tampering types, whereas low-performing models show significant weaknesses, particularly for tampering effects that disrupt temporal coherence (Dropping, Substitution). Moderate-performing models excel slightly in Masking (Figure 8).
Figure 6: Distribution of models into low, moderate, and high-performing categories based on F1 (overall) scores across tampering types.

### v5

@@ Abstract
Multimodal Large Language Models (MLLMs), are recent advancement of Vision-Language Models (VLMs) that have driven major advances in video understanding. However, their vulnerability to adversarial tampering and manipulations remains underexplored. To address this gap, we introduce MVTamperBench, a benchmark that systematically evaluates MLLM robustness against five prevalent tampering techniques: rotation, masking, substitution, repetition, and dropping; based on real-world visual tampering scenarios such as surveillance interference, social media content edits, and misinformation injection. MVTamperBench comprises  3.4K original videos, expanded into over  17K tampered clips covering 19 distinct video manipulation tasks. This benchmark challenges models to detect manipulations in spatial and temporal coherence. We evaluate 45 recent MLLMs from 15+ model families. We reveal substantial variability in resilience across tampering types and show that larger parameter counts do not necessarily guarantee robustness. MVTamperBench sets a new benchmark for developing tamper-resilient MLLM in safety-critical applications, including detecting clickbait, preventing harmful content distribution, and enforcing policies on media platforms. We release all code, data, and benchmark to foster open research in trustworthy video understanding.
@@
Amit Agarwal1*+, Srikant Panda2*, Angeline Charles3*, Hitesh Laxmichand Patel5, Bhargava Kumar4, Priyaranjan Pattnayak6, Taki Hasan Rafi7, Tejaswini Kumar4, Hansa Meghwani1, Karan Gupta5, Dong-Kyu Chae7+ 1Liverpool John Moores University 2Birla Institute of Technology 3Christ University 4Columbia University 5New York University 6University of Washington 7Hanyang University *Equal Contribution +Corresponding Authors Correspondence: amit.pinaki@gmail.com, dongkyu@hanyang.ac.kr
@@ 1 Introduction
Multimodal Large Language Models (MLLMs) have catalyzed significant progress in video understanding, enabling a wide array of applications across domains such as surveillance, healthcare, and autonomous systems. However, their growing integration into high-traffic platforms (e.g., Instagram, Facebook, TikTok) has exposed critical vulnerabilities. Specifically, tampered videos are increasingly exploited to bypass platform policies, disseminate harmful content, and promote clickbait, posing serious challenges for content moderation and policy enforcement Kingra et al. (2023); Times of India (2024).
These threats underscore the urgent need to improve the robustness of MLLMs against real-world manipulations. Although video tampering can involve audio modifications, synthetic speech overlays, or deepfake generation, our benchmark focuses exclusively on visual-only manipulations. This choice is driven by the current limitations of existing models, most of which currently lack comprehensive audiovisual processing capabilities.
Unlike adversarial robustness in static images, an area that has been extensively studied, video tampering introduces unique challenges that arise from the interplay of spatial and temporal dynamics. Common manipulation techniques such as frame dropping, masking, repetition, substitution, and rotation disrupt this coherence, frequently resulting in catastrophic model failures. These methods mirror real-world adversarial tactics: substitution injects objectionable material (e.g., nudity, violence) to circumvent detection; dropping eliminates key surveillance evidence; repetition loops footage to obscure illicit activity; masking occludes critical regions; and rotation induces spatial distortions commonly seen in edited or re-uploaded content.
Existing adversarial and traditional approaches, including black-box Jiang et al. (2019) and cross-modal Wei et al. (2021) attacks, primarily address isolated scenarios rather than systematically evaluating diverse tampering types. Furthermore, existing multimodal benchmarks Pattnayak et al. (2024); Agarwal (2021)—such as MMBench-Video Fang et al. (2024), BLINK Fu et al. (2025), and Video-MME Fu et al. (2024)—focus on multimodal comprehension and temporal reasoning but overlook adversarial robustness. For example, MMBench-Video evaluates cross-modal alignment without adversarial testing; BLINK targets long-form temporal reasoning without addressing tampering impacts; and Video-MME, while effective for vision-language alignment, omits tampering-specific tasks. This leaves a critical gap in systematically assessing how MLLMs withstand real-world manipulations.
To bridge this gap, we introduce MVTamperBench, a benchmark specifically designed to evaluate MLLM robustness against five prevalent tampering techniques. Built from 3.4K original videos—expanded into over 17K tampered clips spanning 19 video tasks—MVTamperBench challenges models to detect manipulations by disrupting temporal and spatial coherence in a video. By focusing on tampering resilience, we believe that MVTamperBench provides a critical tool for improving MLLM robustness, and eventually, it enables the development of tamper-resilient models applicable to real-world challenges such as clickbait detection, content moderation, and policy enforcement, advancing adversarial robustness in video understanding for high-stakes domains.
@@
  • Through experiments on 45 MLLMs across 15+ families, we identify critical vulnerabilities across MLLM families. across MLLM families, without any correlation between model size and performance.
  • Our released code enables researchers to integrate additional datasets and adapt or add new tampering, facilitating domain-specific extensions, and supports reproducibility.
@@ 2 Related Work
The development of benchmarks for MLLMs has significantly advanced the evaluation of image and video understanding tasks. They have covered spatial reasoning, temporal comprehension, object detection, common sense inference, and so on. However, the robustness of MLLMs to adversarial manipulations, particularly video tampering, remains underexplored. This section reviews existing benchmarks, which can be categorized into image-based and video-based evaluations, and analyzes their contributions and shared limitations.
@@ 2.1 Image-based Understanding
Benchmarks like BLINK Fu et al. (2025) and MuirBench Wang et al. (2024a) focus on evaluating static visual reasoning. BLINK tests foundational tasks such as depth estimation, forensic detection, and visual correspondence, which cover a diverse set of challenges for spatial reasoning. Similarly, MuirBench extends this evaluation to multi-image tasks, including action recognition and geographic reasoning, by synthesizing information from diverse sources. While these benchmarks have advanced static image understanding, their reliance on single or multiple still images excludes temporal dynamics and limits their applicability to scenarios involving sequential manipulations, such as those found in video content.
@@ 2.2 Video-based Understanding
Figure 1: Illustration of the five video frame tampering techniques. Each row shows a specific tamper type applied to a video snippet, describing the respective impact on temporal/spatial coherence.
@@
In summary, despite their contributions to advancing MLLM evaluation, existing benchmarks largely focus on performance under ideal conditions and neglect robustness to tampering or adversarial effects. Techniques such as frame substitution, masking, repetition, dropping, and rotation disrupt temporal coherence and pose unique challenges for multimodal models. The absence of systematic evaluations of these tampering techniques highlights a critical gap in ensuring model reliability for real-world applications like forensic analysis, media verification, and misinformation detection.
@@ 3 MVTamperBench
We introduce MVTamperBench, a comprehensive benchmark designed to systematically evaluate the robustness of MLLMs against video tampering techniques. Through our MVTamperBench which introduces diverse manipulations, we aim to broaden the evaluation landscape, enabling a deeper understanding of model strengths and vulnerabilities under adversarial scenarios. Table 3 compares MVTamperBench with existing benchmarks, highlighting its focus areas, strengths, and unique contributions.
In the following subsections, we detail its construction, design choices, and key features. More details can be found in our code [^1] https://amitbcp.github.io/MVTamperBench/, data [^2] https://hf.co/datasets/Srikant86/MVTamperBench and benchmark  [^3] https://github.com/open-compass/VLMEvalKit repositories.
@@ 3.1 Benchmark Construction
To evaluate MLLM robustness under adversarial conditions, we apply distinct tampering methods as shown in Figure 1Dropping, Masking, Substitution, Repetition, and Rotation—to the 3,487 original MVBench Li et al. (2024e) videos (excluding NTU dataset due to licensing), resulting in a total of 17,435 tampered clips. These manipulations target both spatial and temporal coherence, thereby simulating common real-world tampering scenarios such as deliberate frame editing or slicing from unrelated content.
@@ 3.2 Tampering Techniques
Dropping: Removes a 1-second segment for creating temporal discontinuity.
Masking: Overlays a black rectangle on a 1-second segment. It aims to simulate visual data loss.
Rotation: Rotates a 1-second segment by 180 degrees for introducing spatial distortion.
Substitution: Replaces a 1-second segment with a pre-selected clip from another video, in order to disrupt temporal and contextual flow.
@@
The aforementioned effects are applied uniformly across all videos to ensure consistent and comparable evaluation.
@@ 3.3 Design & Implementation
Tampering Duration (1 Second).
We fix tampering to a 1-second segment to align with reports of minimal but impactful real-world tampering, e.g., short edits on social networks or subtle modifications on surveillance feeds Alicia (2024). Our preliminary experiments revealed that using less than 1 second could be overlooked by certain model sampling mechanisms, whereas longer tampering (>3s) often resembled normal scene transitions, reducing adversarial impact.
Tampering Location (Middle).
All manipulations occur at the video’s midpoint to disrupt central content. Our pilot tests showed that tampering near the start or end risked mimicking scene cuts or information loss, which makes detection less indicative of genuine adversarial robustness.
Substitution Source.
For Substitution, the 1-second clip is randomly chosen from a consistent pool of different videos within MVBench. This ensures uniform difficulty across all samples, thus preventing confounds from domain shifts or overly simplistic substitutes.
Modular & Scalable Framework.
Each technique is encapsulated in a reusable class for facilitating custom parameterization (e.g., tampering duration, location, intensity). Our open-source code will integrate seamlessly with VLMEvalKit Duan et al. (2024), thereby promoting reproducible experiments and easy extension to other datasets or tasks.
To inform our final design decisions, we conducted a series of exploratory experiments examining the effects of tampering position and duration. Results from these alternative design configurations are detailed in Appendix A.2.1.
@@ 3.4 Dataset Scope and Statistics
All 3,487 MVBench videos undergo each of the five tampering types, producing 17,435 tampered clips. The five manipulations are uniformly applied to ensure comparability across different MLLM architectures and training regimes. Figure 2 illustrates the distribution of their durations. We can observed that diverse scenarios from short (3-5s) to extended (>20s) sequences are included.
Figure 2: Distribution of video durations. Our dataset spans a broad range of durations, which can reflect varied real-world conditions.
@@ 3.5 Summary
By enforcing consistent parameters (1s duration, midpoint placement) and systematically applying five tampering methods, MVTamperBench offers a controlled yet flexible platform for evaluating tampering resilience. Our design can be easily extended with additional manipulations (e.g., noise injection, partial masking, positions), deepfakes or integrated into domain-specific contexts like surveillance feeds analysis or clickbait detection. We provide additional details on MVTamperBench, including video sources and associated tasks, in Appendix A.1.
@@
4 Experiments
This section outlines the experimental setup, results, and analysis of 45 models evaluated on our proposed benchmark. Our results highlight their strengths, weaknesses, and actionable insights. We provide further overview of the evaluated MLLMs in Appendix A.3.
4.1 Experimental Setups
Evaluation Protocol
Each model is tasked with identifying whether a video has been tampered with or not. For every video, the model is presented with the following structured prompt:
@@ 4.1 Experimental Setups
The dataset comprises both tampered and non-tampered videos. For each tampered video, the corresponding non-tampered video is included to ensure a balanced distribution. Models must select one of the two options (Yes or No), and their predictions are compared against ground truth labels to determine correctness. This task is repeated for all tampering types.
We also compared structured, general, and chain-of-thought prompts (Appendix A.2.2), finding that general and CoT variants led to higher false positives due to limited temporal reasoning in current MLLMs.
Metrics
The primary evaluation metric is the F1 Score, chosen for its ability to balance precision and recall, particularly in scenarios where misclassifications (false positives and false negatives) can significantly impact robustness evaluation. We believe that this metric is the most suitable for our setup, where models must not only detect manipulations but also avoid false positives on non-tampered videos. To capture performance across all tampering types, we compute individual F1 scores for each tampering type. In addition, we also measure F1 (overall) score as the macro average of these individual F1 scores. We thus highlight per-tamper-type strengths and weaknesses as well as overall model performance.
@@
Figure 5: Scatter plot showing no correlation between model size and overall F1 (Pearson r=0.05).
@@ 4.2 Results and Analysis
We evaluate the performance of 45 MLLMs using F1 (overall) scores and individual F1 scores across five tampering types. The distribution of F1 (overall) scores (Figure 4) reveals significant variability in the robustness of the model, with several models struggling to detect tampering (F1 < 0.2), while a few high-performing models achieve F1 > 0.8.
We did not observe a correlation between model size and tampering detection performance (Pearson correlation = 0.05, Figure 5). This highlights that architectural differences and training techniques, rather than parameter count, contribute more significantly to robustness against tampering.
In addition, Figure 3 showcases model-specific adaptability across different tampering types. Models with F1 (overall) > 0.8 are consistently robust across all tampering types, while those with F1 < 0.2 perform slightly better on Masking, which relies more on spatial reasoning. Models in between generally struggle with temporal disruptions like Dropping and Substitution, reflecting challenges in temporal coherence.
4.2.1 Analysis based on Performance
Figure 6 categorizes MLLMs into low-, moderate-, and high- performing groups based on their F1-score distribution. We define the boundaries using the 0.25 quantile (F1 = 0.071) and 0.75 quantile (F1 = 0.846). We round-off the quantile thresholds to the nearest integers to identify low-performing models (F1 (overall) < 0.01) and high-performing models (F1 (overall) > 0.8).
Figure 7 illustrates the average F1 scores for each category. We observe stark differences between groups: High-performing models maintain consistent performance across all tampering types, whereas low-performing models show significant weaknesses, particularly for tampering effects that disrupt temporal coherence (Dropping, Substitution). Moderate-performing models excel slightly in Masking (Figure 8).
Figure 6: Distribution of Number of Models in low, moderate, and high-performing categories based on F1 (overall).